# Generalized quantum Zernike Hamiltonians: Polynomial Higgs-type algebras and algebraic derivation of the spectrum

### v1

@@ 6 Higher-order quantum Zernike Hamiltonians: Conjectures
Second, to ensure that both spectra $E_{\rm I}$ and $E_{\rm II}$ are real, it is necessary to require that $\gamma_{k}$ is purely imaginary for odd values of $k$, while it must be real for even values of $k$. Third, the above results are well-defined for any value of all the coefficients $\gamma_{k}$. In particular, we can take a single or more coefficients different from zero. For instance, the simplest choice is to set only $\gamma_{1}\neq 0$, leading to a linear spectrum in $n$, that corresponds to the first system with a discrete spectrum; in fact, it has been interpreted as the isotropic Euclidean oscillator in section 3.2 (when $\gamma_{k}=0$ for all $k$, we recover the trivial geodesic motion on the Euclidean plane). Finally, from this point of view, the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (2.3) can be regarded as the superposition of $N$ possible potentials, each of them determined by one $\gamma_{k}\neq 0$, leading to a term of $k^{{\rm th}}$-order in the momenta in the quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$, as well as a term $n^{k}$ (or $(n+2)^{k}$) in the spectrum.
We now outline the main results for the case $N=5$, which are in full agreement with both conjectures. The quantum symmetry ${\hat{I}}^{\prime}_{5}$ in (2.6) of ^$\hat{\mathcal{H}}_{5}$ is explicitly obtained, after some cumbersome computations, as
^$\hat{\mathcal{H}}_{5}={\hat{I}}_{5}+{\hat{I}}_{5}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
Next we introduce the operators $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8), from which the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ with the same expression as in (5.1) can be defined. These operators close a polynomial symmetry algebra $\mathfrak{sl}^{(9)}(2,\mathbb{R})$ fulfilling the commutation rules (2.9), in such a manner that the factorizing terms in (2.10) are given by
Type $\begin{array}[]{lll}\mbox{\rm Type I:}&\displaystyle u_{\rm I}=-\frac{{n}}{2},&\quad\displaystyle E_{\rm I}=-\bigl{(}{\rm i}\gamma_{1}{n}+\gamma_{2}{n}^{2}-{\rm i}\gamma_{3}{n}^{3}-\gamma_{4}{n}^{4}+{\rm i}\gamma_{5}n^{5}\bigr{)}.\\[10.0pt] \mbox{\rm Type II:}&\displaystyle u_{\rm II}=-\frac{{n}}{2},&\quad E_{\rm II}={\rm i}\gamma_{1}({n}+2)-\gamma_{2}({n}+2)^{2}-{\rm i}\gamma_{3}({n}+2)^{3}+\gamma_{4}({n}+2)^{4}+{\rm i}\gamma_{5}(n+2)^{5}.\end{array}$
These formulae coincide with those obtained from both conjectures when substituting the value of $N$.
In summary, in order to consistently prove the conjectures 1 and 2, it is first necessary to find the generic expression of the quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ in (2.6), from which the structure of the polynomial Higgs-type algebra of order $(2N-1)^{{\rm th}}$ is obtained, which constitutes the main computational obstruction. A second step consists in finding the precise structure of the quantum operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$, starting from $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8), leading to the factorization terms in (6.1). As shown above, this is far from being a trivial task, even with the use of symbolic computer packages.
@@ 7 Conclusions and outlook
In this work, a quantum version of the generalized classical Zernike Hamiltonian first studied in [3] has been proposed, showing that one of its main features, namely the superintegrability property, is preserved by quantization. In addition, it has been shown that the spectrum of the quantum Hamiltonian (2.3) can be inferred by purely algebraic means, using the quantum symmetries and the structure of the associated polynomial Higgs-type algebra. Although explicit computations become extremely cumbersome for values $N\geq 6$, a generic prescription has been given to solve the spectrum problem for any given value of $N$. Based on the general shape of the factorizing terms of the structure function and the resulting spectra for the admissible solutions that allow us to interpret the result as a higher-order perturbation of the usual quantum Zernike model, we conjecture the solutions for arbitrary value of $N$, without making explicit use of the quantum symmetries, for which the exact expression is still an open problem. In this context, another unanswered question concerns the number of solutions of equation (2.14), and whether for higher values of $N$ we can find additional solutions that are well-defined and behave properly for the limits $\gamma_{k}\rightarrow 0$ for $k\geq 3$, albeit this seems not probable in view of the structure observed for values $N\leq 5$.
Furthermore, we stress that the interpretation of the Hamiltonian $\mathcal{H}_{2}$ (2.3) as defining, in a unified form, the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$, have allowed us to consider the generalized Zernike systems as superintegrable perturbations for these oscillators when $N\geq 3$. Moreover, we have also shown that such perturbations can be of spherical or hyperbolic type depending on the value of the corresponding coefficient $\gamma_{k}$. In this sense, these new results extend Bertrand’s theorem to the momentum-dependent systems studied here.
@@
  • The Schrödinger equation could be faced for the simplest flat Euclidean system at least for $N=3$ and the resulting wave eigenfunctions and spectrum analyzed according to the spherical or hyperbolic type of the perturbation. This result can provide the path to solve the Schrödinger equation for the perturbations of curved oscillators on $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$.
  • The generalized quantum Zernike Hamiltonians could be implemented in the three Lorentzian spaces of constant curvature, Minkowski, anti-de Sitter and de Sitter spacetimes, by means of graded contractions or analytical continuations [40]; the appropriate coordinates would be the geodesic polar ones shown in (3.9).
@@
Let us consider the Lie algebra $\mathfrak{gl}(2)=\operatorname{span}\{J_{-},J_{+},J_{3},\Xi\}$ with commutation relations and Casimir operator ${C}$ given by
$[J_{3},J_{+}]=2{\rm i}J_{+},\qquad[J_{3},J_{-}]=-2{\rm i}J_{-},\qquad[J_{-},J_{+}]=4{\rm i}J_{3}+2\,\Xi,\qquad[\Xi\,,\,\cdot\,]=0,$ (7.1)
@@
By construction [48, 49], $\hat{C}^{(2)}$ commutes with the operators (7.4), so that any smooth function ${H}$ defined on them determines a two-dimensional quantum integrable Hamiltonian:
@@
The coassociative property of the coproduct allows one to extend this result to arbitrary dimension $d$ providing $(2d-3)$ ‘universal’ algebraically independent operators [49] for $\hat{H}^{(d)}$, so that the resulting systems are called quasi-maximally superintegrable since only one additional quantum symmetry is needed to ensure maximal superintegrability. In the classical case, the central generator $\Xi$ vanishes, so that $(\mathfrak{gl}(2),\Delta)$ reduces to a Poisson $\mathfrak{sl}(2,\mathbb{R})$-coalgebra, which has been deeply studied [50, 49]. Recall also that the Racah algebra for $\mathfrak{sl}(2,\mathbb{R})$ has been widely studied in [51, 52, 53].
@@
Therefore, if we set $\lambda_{1}=\lambda_{2}=0$ we recover the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3), while the operator $\hat{C}^{(2)}$ in (7.5) gives the square of the quantum angular momentum operator (2.5). Hence, for arbitrary coefficients $\lambda_{1}$ and $\lambda_{2}$, the Hamiltonian $\hat{H}^{(2)}$ is formed by the superposition of ^$\hat{\mathcal{H}}_{N}$ with two Rosochatius–Winternitz potentials. In the classical case with $N=2$ and within the interpretation of ^$\hat{\mathcal{H}}_{2}$ as the Hamiltonian of a curved oscillator, $\hat{H}^{(2)}$ is just the curved Smorodinsky–Winternitz system which is known to be superintegrable [50]; the presence of a positive $\lambda_{i}$-potential corresponds to introduce a ‘centrifugal’ barrier on $\mathbf{E}^{2}$ and $\mathbf{H}^{2}$, but a noncentral oscillator on $\mathbf{S}^{2}$ [36]. From this perspective, it is rather natural to look for an additional quantum symmetry for $\hat{H}^{(2)}$ (7.6) and to deduce and analyze the resulting spectrum by applying an algebraic approach.
@@ Acknowledgements
R.C.-S., F.J.H. and A.B. have been partially supported by Agencia Estatal de Investigación (Spain) under the grant PID2023-148373NB-I00 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. F.J.H. and A.B. also acknowledge support by the Q-CAYLE Project funded by the Regional Government of Castilla y León (Junta de Castilla y León, Spain) and by the Spanish Ministry of Science and Innovation (MCIN) through the European Union funds NextGenerationEU (PRTR C17.I1). The research of D.L. has been partially funded by MUR - Dipartimento di Eccellenza 2023-2027, codice CUP G43C22004580005 - codice progetto DECC23__\__012__\__DIP and partially supported by INFN-CSN4 (Commissione Scientifica Nazionale 4 - Fisica Teorica), MMNLP project. D.L. is a member of GNFM, INdAM. I.M. has been supported by Australian Research Council Future Fellowship FT180100099.

### v2

@@ 5 Higher-order quantum Zernike Hamiltonians: Conjectures
Second, to ensure that both spectra $E_{\rm I}$ and $E_{\rm II}$ are real, it is necessary to require that $\gamma_{k}$ is purely imaginary for odd values of $k$, while it must be real for even values of $k$. Third, the above results are well-defined for any value of all the coefficients $\gamma_{k}$. In particular, we can take a single or more coefficients different from zero. For instance, the simplest choice is to set only $\gamma_{1}\neq 0$, i.e., $N=1$, leading to a linear spectrum in $n$, that corresponds to the first system with a discrete spectrum; in fact, it has been interpreted as the isotropic Euclidean oscillator in section 3.2 (when $\gamma_{k}=0$ for all $k$, we recover the trivial geodesic motion on the Euclidean plane). Finally, from this point of view, the generalized quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{N}$ in (2.3) can be regarded as the superposition of $N$ possible potentials, each of them determined by one parameter $\gamma_{k}\neq 0$, leading to a term of $k^{{\rm th}}$-order in the momenta in the quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$, as well as a term $n^{k}$ (or $(n+2)^{k}$) in the spectrum.
[…]
@@ 6 Conclusions and outlook
In this work, a quantum version of the generalized classical Zernike Hamiltonian, first studied in [4], has been proposed, showing that one of its main features, namely the superintegrability property, is preserved by quantization. In addition, it has been shown that the spectrum of the quantum Hamiltonian (2.3) can be inferred by purely algebraic means, using the quantum symmetries and the structure of the associated polynomial Higgs-type algebra. Although explicit computations become extremely cumbersome for values $N\geq 6$, a generic prescription has been given to solve the spectrum problem for any given value of $N$. Based on the general shape of the factorizing terms of the structure function and the resulting spectra for the admissible solutions that allow us to interpret the result as a higher-order perturbation of the usual quantum Zernike model, we conjecture the solutions for arbitrary value of $N$, without making explicit use of the quantum symmetries, for which the exact expression is still an open problem. In this context, another unanswered question concerns the number of solutions of equation (2.14), and whether for higher values of $N$ we can find additional solutions that are well-defined and behave properly for the limits $\gamma_{k}\rightarrow 0$ for $k\geq 3$, albeit this seems not probable in view of the structure observed for values $N\leq 5$.
Furthermore, we stress that the interpretation of the Hamiltonian $\mathcal{H}_{2}$ (2.3) as defining, in a unified form, the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$, have allowed us to consider the generalized Zernike systems as superintegrable perturbations for these oscillators when $N\geq 3$. Moreover, we have also shown that such perturbations can be of spherical or hyperbolic type, depending on the value of the corresponding coefficient $\gamma_{k}$. In this sense, these new results extend Bertrand’s theorem to the momentum-dependent systems studied here.
@@
  • The Schrödinger equation could be faced for the simplest flat Euclidean system, corresponding to $N=1$, at least for the cubic $N=3$ perturbation, and the resulting wave eigenfunctions and spectrum analyzed according to the spherical or hyperbolic type of the perturbation. This result can provide the path to solve the Schrödinger equation for the perturbations of curved oscillators on $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$.
  • The generalized quantum Zernike Hamiltonians could be implemented in the three Lorentzian spaces of constant curvature, Minkowski, anti-de Sitter and de Sitter spacetimes, by means of graded contractions or analytical continuation [44]; the appropriate coordinates would be the geodesic polar ones shown in (3.9).
@@
Let us consider the Lie algebra $\mathfrak{gl}(2)=\operatorname{span}\{J_{-},J_{+},J_{3},\Xi\}$ with commutation relations and Casimir operator C{C} given by
$J_{3},J_{+}]=2{\rm i}J_{+},\qquad[J_{3},J_{-}]=-2{\rm i}J_{-},\qquad[J_{-},J_{+}]=4{\rm i}J_{3}+2\,\Xi,\qquad[\Xi\,,\,\cdot\,]=0,$ (6.1)
@@
By construction [52, 53], $\hat{C}^{(2)}$ commutes with the operators (6.4), so that any smooth function H{H} defined on them determines a two-dimensional quantum integrable Hamiltonian:
@@
The coassociative property of the coproduct allows one to extend this result to arbitrary dimension $d$ providing (2d3)(2d-3) ‘universal’ algebraically independent operators [53] for $\hat{H}^{(d)}$, so that the resulting systems are called quasi-maximally superintegrable since only one additional quantum symmetry is needed to ensure maximal superintegrability. In the classical case, the central generator $\Xi$ vanishes, so that $(\mathfrak{gl}(2),\Delta)$ reduces to a Poisson $\mathfrak{sl}(2,\mathbb{R})$-coalgebra, which has been deeply studied [54, 53]. Recall also that the Racah algebra for $\mathfrak{sl}(2,\mathbb{R})$ has been widely studied in [55, 56, 57].
@@
Therefore, if we set $\lambda_{1}=\lambda_{2}=0$ we recover the generalized quantum Zernike Hamiltonian $\hat{\mathcal{H}}_{N}$ (2.3), while the operator $\hat{C}^{(2)}$ in (6.5) gives the square of the quantum angular momentum operator (2.5). Hence, for arbitrary coefficients $\lambda_{1}$ and $\lambda_{2}$, the Hamiltonian $\hat{H}^{(2)}$ is formed by the superposition of $\hat{\mathcal{H}}_{N}$ with two Rosochatius–Winternitz potentials. In the classical case with $N=2$, and within the interpretation of $\hat{\mathcal{H}}_{2}$ as the Hamiltonian of a curved oscillator, $\hat{H}^{(2)}$ is just the curved Smorodinsky–Winternitz system which is known to be superintegrable [54]; the presence of a positive $\lambda_{i}$-potential corresponds to introduce a ‘centrifugal’ barrier on $\mathbf{E}^{2}$ and $\mathbf{H}^{2}$, but a noncentral oscillator on $\mathbf{S}^{2}$ [40]. From this perspective, it is rather natural to look for an additional quantum symmetry for $\hat{H}^{(2)}$ (6.6) and to deduce and analyze the resulting spectrum by applying an algebraic approach.
@@ Acknowledgements
R.C.-S., F.J.H. and A.B. have been partially supported by Agencia Estatal de Investigación (Spain) under the grant PID2023-148373NB-I00 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. F.J.H. and A.B. also acknowledge support by the Q-CAYLE Project funded by the Regional Government of Castilla y León (Junta de Castilla y León, Spain) and by the Spanish Ministry of Science and Innovation (MCIN) through the European Union funds NextGenerationEU (PRTR C17.I1). R.C.-S. also acknowledges support by the PID2024-156578NB-I00 funded by MICIU /AEI /10.13039/501100011033 / FEDER, EU. The research of D.L. has been partially funded by MUR - Dipartimento di Eccellenza 2023-2027, codice CUP G43C22004580005 - codice progetto DECC23_\_012_\_DIP and partially supported by INFN-CSN4 (Commissione Scientifica Nazionale 4 - Fisica Teorica), MMNLP project. D.L. is a member of GNFM, INdAM. I.M. has been supported by Australian Research Council Future Fellowship FT180100099.
Appendix A The fourth-order quantum Zernike Hamiltonian
We proceed similarly to sections 3 and 4, first giving the quantum symmetries and the possible spectra for the quartic Hamiltonian $\hat{\mathcal{H}}_{4}$ in (2.3), depending on four arbitrary parameters $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$, and imposing that the quantum quadratic Hamiltonian $\hat{\mathcal{H}}_{2}$ in (2.3) is recovered for the vanishing of $\gamma_{3}$ and $\gamma_{4}$.
The two quantum symmetries (2.6) of $\hat{\mathcal{H}}_{4}$, now of fourth-order in the momenta, read
and verify the algebraic dependence relation
$\hat{\mathcal{H}}_{4}={\hat{I}}_{4}+{\hat{I}}_{4}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
[…]
$\begin{split}\hat{\mathcal{K}}&:=\frac{1}{2}\hat{\mathcal{K}}_{1},\\ \hat{\mathcal{K}}_{+}&:=\hat{\mathcal{K}}_{2}+\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl(\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr)\hat{\mathcal{K}}_{1}^{2},\\ \hat{\mathcal{K}}_{-}&:=\hat{\mathcal{K}}_{2}-\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl(\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr)\hat{\mathcal{K}}_{1}^{2},\end{split}$ (A.1)
and close on the commutation relations in (2.9) with
which further factorizes as
$\begin{split}\Phi(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\qquad\bigl[\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\bigr]=0,\\ \Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\frac{1}{4}\bigl(\hat{\mathcal{H}}_{4}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}+8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\bigr),\\[2.0pt] \Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\hat{\mathcal{H}}_{4}-2({\rm i}\gamma_{1}-2\gamma_{2}-4{\rm i}\gamma_{3}+8\gamma_{4}){\rm Id}+2({\rm i}\gamma_{1}-4\gamma_{2}-12{\rm i}\gamma_{3}+32\gamma_{4})\hat{\mathcal{K}}\\[2.0pt] &\qquad+4(\gamma_{2}+6{\rm i}\gamma_{3}-24\gamma_{4})\hat{\mathcal{K}}^{2}-8({\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\\[2.0pt] &=\hat{\mathcal{H}}_{4}+2{\rm i}\gamma_{1}\bigl(\hat{\mathcal{K}}-{\rm Id}\bigr)+4\gamma_{2}\bigl(\hat{\mathcal{K}}-{\rm Id}\bigr)^{2}-8{\rm i}\gamma_{3}\bigl(\hat{\mathcal{K}}-{\rm Id}\bigr)^{3}-16\gamma_{4}\bigl(\hat{\mathcal{K}}-{\rm Id}\bigr)^{4}.\end{split}$ (A.2)
We introduce the finite-dimensional representation of (2.12) into the expressions (A.2), obtaining the structure function $\Phi=\Phi_{1}\Phi_{2}$ (2.13) as
$\begin{split}\Phi_{1}({B},E,u)&=\frac{1}{4}\bigl(E-2{\rm i}\gamma_{1}({B}+u)+4\gamma_{2}({B}+u)^{2}+8{\rm i}\gamma_{3}({B}+u)^{3}-16\gamma_{4}({B}+u)^{4}\bigr),\\[2.0pt] \Phi_{2}({B},E,u)&=E+2{\rm i}\gamma_{1}({B}+u-1)+4\gamma_{2}({B}+u-1)^{2}-8{\rm i}\gamma_{3}({B}+u-1)^{3}-16\gamma_{4}({B}+u-1)^{4}.\end{split}$ (A.3)
The spectrum $E$ of $\hat{\mathcal{H}}_{4}$ (2.4) is derived by imposing the two conditions (2.14) and is established as follows.
Proposition 3
(i) The set of equations (2.14) obtained from (A.3) leads to the corresponding solutions for the representation dependent constant u=u(n)u=u({n}) and the spectrum E=E(n)E=E({n}) of the Hamiltonian $\hat{\mathcal{H}}_{4}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They provide the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ through (A.3).
(ii) Among all the solutions, there are only two types such that the limits $\gamma_{4}\to 0$ and $\gamma_{3}\to 0$ are simultaneously well-defined for any value of $\gamma_{1}$ and $\gamma_{2}$, namely
[…]
The excluded solutions exhibit divergencies, similarly to the solution (4.2) for the cubic Hamiltonian $\hat{\mathcal{H}}_{3}$ in section 4, although the resulting expressions are more complicated in this case (they involve both parameters $\gamma_{3}$ and $\gamma_{4}$ within the denominators). In this sense, the discarded solutions in propositions 2 and 3 can be regarded as ‘spurious’.
It is worth remarking that the results given in proposition 3 are well-defined for any value of the four parameters $\gamma_{k}$ $(k=1,\dots,4)$. As expected, they lead exactly to the results presented in proposition 2 for $\gamma_{4}=0$ and, then, to those characterized by proposition 1 when $\gamma_{3}=0$. In addition, it is possible to set $\gamma_{3}=0$ in the solutions (A.4), keeping the parameter $\gamma_{4}$, which determines a superintegrable quartic perturbation of the initial Hamiltonian $\hat{\mathcal{H}}_{2}$.
We observe that the expressions (A.4) take real values whenever the odd coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while the even ones $\gamma_{2}$ and $\gamma_{4}$ take real values. This difference between odd/even parameters has already been studied in the classical picture [4]. Therefore, by taking into account (4.3), let us denote
obtaining two types of spectra for $\hat{\mathcal{H}}_{4}$ given by
to be compared with (4.4). In addition, if we set $\alpha=-1$ and $\beta=-2$ and interpret n{n} as the principal quantum number $\mathfrak{n}$, we find two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian $\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) depending on two real parameters $\mu$ and $\nu$, and given by
Appendix B The fifth-order quantum Zernike Hamiltonian
We now outline the main results for the case $N=5$, which are in full agreement with both conjectures proposed in section 5. The quantum symmetry ${\hat{I}}^{\prime}_{5}$ in (2.6) of $\hat{\mathcal{H}}_{5}$ is explicitly obtained, after some cumbersome computations, as
$\hat{\mathcal{H}}_{5}={\hat{I}}_{5}+{\hat{I}}_{5}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
[…]
$\begin{array}[]{lll}\mbox{\rm Type I:}&\displaystyle u_{\rm I}=-\frac{{n}}{2},&\quad\displaystyle E_{\rm I}=-\bigl({\rm i}\gamma_{1}{n}+\gamma_{2}{n}^{2}-{\rm i}\gamma_{3}{n}^{3}-\gamma_{4}{n}^{4}+{\rm i}\gamma_{5}n^{5}\bigr).\\[10.0pt] \mbox{\rm Type II:}&\displaystyle u_{\rm II}=-\frac{{n}}{2},&\quad E_{\rm II}={\rm i}\gamma_{1}({n}+2)-\gamma_{2}({n}+2)^{2}-{\rm i}\gamma_{3}({n}+2)^{3}+\gamma_{4}({n}+2)^{4}+{\rm i}\gamma_{5}(n+2)^{5}.\end{array}$
It can be directly verified that all of the above expressions do coincide with those obtained from both conjectures in section 5, by entering the value $N=5$.